OpenStudy (anonymous):

Drew is making a cardboard cut out of a right triangle. He knows that the longest side must be 15 inches long. One of the short sides must be 8 inches. How many inches long will he make the other side? Use the appropriate formula to find the solution and round to the nearest tenth. Show all steps or reasoning used. Be sure to check your answer for reasonableness. YOU GUYS PLEASE HELP ME ILL RETURN THE FAVORR

OpenStudy (anonymous):

The longest side of a right triangle is the hypotenuse. Relevent equation are the pythagorean theorem : a^2 + b^2 = c^2 15^2= 8^2 + x^2 x=12.7 inches

OpenStudy (anonymous):

thank you this really helped me out :)